Single proportion test

According to an October 2019 study by the American Association of Endodontists (AAE) of 1000 Americans, 53% of survey respondents said that they would rather have a snake in their lap for 15 minutes than get a root canal. You are skeptical of this because snakes are creepy. If you take a survey of 306 people and find that 162 would rather have a snake in their lap for 15 minutes than get a root canal, what is the value of your test statistic? (z subscript t e s t end subscript)

Round your answer to two (2) decimal places
Answer #1

Solution :

\hat p = x / n = 162 / 306 = 0.5294

P0 = 0.53

1 - P0 = 1-0.53 = 0.47

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

=0.5294 - 0.53 / [\sqrt0.53*(0.47) /306 ]

= -0.02